write an equation of the line that passes through (5, -4) and is parallel to y=2x+3
Found 2 solutions by math_helper, MathTherapy:
Answer by math_helper(2461)


y = 2x+3 has the form y = mx + b (m=slope, b=y-intercept).
Parallel lines will have the equal slopes.
So we know this much: the parallel line has the form: y = 2x + b
You can either use the point (5,-4) to solve for b, or you
can use the point-slope form of a line y-y0 = m(x-x0) and
plug in (5,-4) for (x0,y0).
Solving for b:
-4 = 2(5) + b --> b = -14 ==> eqn is y = 2x - 14
Using point-slope form:
y - (-4) = 2(x - 5)
y + 4 = 2x - 10
y = 2x - 14 (same result)

Answer by MathTherapy(10552)


write an equation of the line that passes through (5, -4) and is parallel to y=2x+3
y = 2x + 3, in standard form: 2x - y = - 3
Equation requested: